3. The Firms Problem and Taxes (30 points total). Suppose the representative firm has a production function given by Y = 2K(b) (5 points) Write down the firms optimality condition. (Hint: Be careful when computing the marginal product of labor for
Answer #1

A) nominal profit= p*Y-p*real wage* N - p*Y*t

Real profit=y - real wage -yt

Profit=8√N -2N-8*t*√N

Profit=8√N(1-t) - 2N

B) optimal condition of firm,

MPL*1-t=real wage

MPL=4/√N

(4-4t)/√N=2

C) labour demand;(4-4t)/√N=2

√N=2-2t

N=(2-2t)^2

N=4-4t2-8t

Output;y=8√N=8√(2-2t)^2=8(2-2t)=16-16t

Profit=16-16t(1-t)-2(4-4t2-8t)=16-16t+16t2-8+8t2+16t

Profit=8-8t2

Tax Increase, decreases labour demand .

D)profit Maximizing problem;,

Gross Profit=8√N-2N

Net profit( after tax)=8√N*t- 2*N*t

Optimal condition,( now optimal condition remain same as without tax because tax is levied on profit,so objective is Maximizing profit)

MPL=real wage

4/√N=2

2=√N

N=4